OpenStudy (anonymous):

Factor: x2 - 9x + 12

OpenStudy (anonymous):

Did you type the problem wrong? If the 9 was 8... then (x-6)(x-2)

OpenStudy (anonymous):

You will need to use the quadratic equation or complete the square to factor this quadratic equation. \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] In your equation, b=-9 a=1 and c=12 \[(9\pm \sqrt{(-9)^{2}-[(4)(1)(12)]})/[(2)(1)]\] Which becomes \[9/2\pm (\sqrt{81-48})/2\] Which becomes \[9/2\pm(\sqrt{33})/2\] So your factors are \[9/2+\sqrt{33}/2\]\[9/2-\sqrt{33}/2\]